# Langevin equation with potential of mean force: The case of anchored bath

### v1

@@ VI Klein-Gordon chain: the PMF
Let us now consider a more specific model where the bath is formed by the Klein-Gordon harmonic chain. The chain consists of $N\gg 1$ atoms of mass $m$, connected by springs with the stiffness $k_{0}$. Each atom of the chain is subjected to the on-site anchor harmonic potential with the stiffness $k_{a}$. The system of interest is a particle of mass $M$, connected by a $k_{0}$-spring to the first atom of the chain and subjected to an external potential $V_{ex}(X)$. We assume a boundary condition when the last particle of the chain is connected by a $k_{0}$-spring to a wall. The model is sketched in Fig. 1 and can be viewed as the familiar Rubin model Zwanzig ; Weiss , extended by the presence of the anchoring potentials acting on atoms of the bath and the external potential $V_{ex}(X)$ acting on the system. It is well known that a non-anchored chain can serve, in the limit $N\to\infty$, as a passive bath Zwanzig ; Weiss . In contrast, the Klein-Gordon chain is non-passive: in thermal equilibrium it produces a systematic harmonic potential acting on the system, so the PMF has the form
@@
Below we evaluate the strength $\kappa$ of the bath-induced potential; the result is given by Eq. (113).
@@
Figure 1: The setup considered in the paper. The black circle represents the system of interest, which is a classical particle subjected to an (arbitrary) external potential $V_{ex}(X)$. The system is connected by the spring with stiffness $k_{0}$ to the thermal bath, formed by the Klein-Gordon lattice. The latter consists of $N$ particles connected by the springs with stiffness $k_{0}$. Each particle of the lattice is subjected to on-site (anchoring) harmonic potentials with the stiffness $k_{a}$.
@@
The transformation has the same form as for the corresponding harmonic chain with no anchoring on-site potential. It is canonical and preserves phase-space volumes, i. e. the Jacobian of the transformation equals $1$. Therefore, the average $\langle\cdots\rangle$ over the bath variables $\{x_{i},p_{i}\}$, which we used in the previous sections, can be replaced by the average over $\{Q_{j},P_{j}\}$.
@@
$H_{0}=\frac{1}{2}\sum_{j}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right]^{2}\right\}+\frac{1}{2}\,\kappa\,X^{2},$ (112)
@@
Thus, the bath formed by the infinite Klein-Gordon chain is non-passive; it acts with the systematic average force $\langle F\rangle=-\kappa\,X$ on the system fixed at $X$. The effective stiffness $\kappa$ of that force as a function of $a=\omega_{a}/\omega_{0}$ is given by Eq. (113). It monotonically increases from 0 at $a=0$ to $k_{0}$ for $a\gg 1$. Both limiting values are intuitively plausible. For $a=0$, the model is equivalent to the Rubin model Zwanzig ; Weiss where the chain is not anchored and in the limit $N\gg 1$ (but not for finite $N$) serves as a passive bath, which does not produce a systematic force on the fixed system. For $a\gg 1$ the chain acts as an infinitely heavy anchor. In that case, the system feels only the spring of stiffness $k_{0}$ by which it is attached to the terminal atom of the chain.
@@ VII Klein-Gordon chain: Langevin equation
$F=-\frac{\partial H_{0}^{\prime\prime}}{\partial X}=-k_{0}X+\sum_{j}c_{j}Q_{j},$ (119)
where $c_{j}$ are given by Eq. (106). Then $F_{0}(t)=e^{L_{0}t}F$ is
$F_{0}(t)=-k_{0}X+\sum_{j}c_{j}Q_{j}(t),$ (120)
where the dynamics of $Q_{j}(t)$ is governed by Hamiltonian $H_{0}$, Eq. (109), and the equations of motion
@@
We observe that $F_{0}(t)$ satisfies conditions (92)-(94) mentioned at the end of Section 5 and tehrefore statistical properties of the zero-centered noise $E_{0}(t)=F_{0}(t)-\langle F\rangle$ do not depend on $X$. We can also show that directly. Indeed, using the substitution
@@
the spring constant $\kappa$ of the bath-induced potential is given by Eq. (113), and the dissipative kernel $\Gamma(t)$ is related to the zero-centered noise $E_{0}(t)$ by the standard fluctuation-dissipation relation
@@
and restricting to the leading approximation $F^{\dagger}(t)\approx E_{0}(t)$. However, one can show that for the presented model the next term $E_{1}(t)$, which is given by Eq. (83), vanishes. Indeed, as follows from Eqs. (125) and (129), $E_{0}(t)$ is linear in $X$, the derivative
@@
We finish this section by deriving a more explicit expression for the correlation $\langle EE_{0}(t)\rangle$ and the dissipation kernel $K(t)$. According to Eqs. (130) and (133), the kernel equals
@@
and the noise correlation equals $\langle EE_{0}(t)\rangle=\frac{M}{\beta}\,K(t)$. For the non-anchored chain, that is, for $a=0$, the dimensionless kernel $\gamma(t)$ is expressed in terms of the Bessel function $J_{1}(x)$,
$\gamma(t)=\frac{1}{2\,t}\,J_{1}(\omega_{0}t).$ (140)
@@ VIII Conclusion
The concept of the PMF naturally emerges in equilibrium statistical mechanics of complex systems consisting of many particles. The force $F_{i}$ on $i$-th particle of the system is produced by the bath disturbed by the interaction with other parts of the system. As a result, even if the bath is a translationally-invariant fluid, the mean force $\langle F_{i}\rangle$ is generally not zero and is determined by the bath-induced part of the PMF, denoted above as $\Delta V$. Although for equilibrium systems the concept is straightforward, there is a variety of views on how to incorporate the PMF in the Langevin description of systems out of equilibrium.
In this paper, we have simplified the problem by assuming that the system of interest consists of a single particle. In that case, the PMF is usually irrelevant because the homogeneous bath is passive and produces no systematic force on a fixed system. That is not so, however, if the bath is not translationally-invariant. In the presented model the symmetry of the bath is broken by on-site anchor potentials applied to individual molecules of the bath. Although that setting offers a simpler exposition, we were unable to derive a Langevin equation in a closed form. The non-passivity of the bath not only modifies the average potential acting on the system, but also makes the dissipation kernel and noise statistics depend on the position of the system $X$. That dependence is a priory unknown and there seems to be no general relation connecting it to the PMF. Therefore, we argue that there is no general form for a Langevin equation with the PMF. The conclusion corroborates the findings of Refs. Schilling ; Schilling_review ; Vroy , although we used here a different projection operator technique.
@@
$\dot{P}(t)=-\frac{\partial V_{ex}}{\partial X}-\kappa\,X(t)-\int_{0}^{t}\Gamma(t-\tau)\,P(\tau)\,d\tau+\zeta(t)-M\,X\,K(t),$ (A8)
@@
$K(t)=\frac{1}{M}\,\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\cos(\Omega_{j}t),\quad\zeta(t)=\sum_{j=1}^{N}c_{j}Q_{j}^{0}(t),$ (A9)
@@
and $X=X(0)$. In the limit $N\to\infty$ $\kappa$ is given by expression (113). Finally, absorbing the term MXK(t)-M\,X\,K(t) into noise and using notations of the main text, we obtain

### v2

@@ VI Klein-Gordon chain: the PMF
Let us now consider a more specific model where the bath is formed by the Klein-Gordon harmonic chain. The chain consists of $N\gg 1$ atoms of mass $m$, connected by springs with stiffness $k_{0}$. Each atom of the chain is subjected to the on-site anchor harmonic potential with stiffness $k_{a}$. The system of interest is a particle of mass $M$, connected by a $k_{0}$-spring to the first atom of the chain and subjected to external potential $V_{ex}(X)$. We assume a boundary condition when the terminal atom of the chain is connected by a $k_{0}$-spring to a wall. The model is sketched in Fig. 1 and can be viewed as the familiar Rubin model Zwanzig ; Weiss , extended by the presence of anchor potentials acting on atoms of the bath and external potential $V_{ex}(X)$ acting on the system. It is well known that a non-anchored chain can serve, in the limit $N\to\infty$, as a passive bath Zwanzig ; Weiss . In contrast, the Klein-Gordon chain is non-passive: in thermal equilibrium it produces a systematic harmonic potential acting on the system, so the PMF has the form
@@
Below we evaluate stiffness $\kappa$ of the bath-induced potential; the result is given by Eq. (110).
@@
Figure 1: The setup considered in Sections 6 and 7. The black circle represents the system of interest, which is a classical particle subjected to an (arbitrary) external potential $V_{ex}(X)$. The system is connected by the spring with stiffness $k_{0}$ to the thermal bath, formed by the Klein-Gordon lattice. The latter consists of $N$ atoms connected by the springs with stiffness $k_{0}$. Each atom of the lattice is subjected to on-site (anchoring) harmonic potentials with stiffness $k_{a}$.
@@
The transformation has the same form as for the corresponding harmonic chain with no anchor potentials. It is canonical and preserves phase-space volumes, i. e. the Jacobian of the transformation equals $1$. Therefore, the average $\langle\cdots\rangle$ over the bath variables $\{x_{i},p_{i}\}$, which we used in the previous sections, can be replaced by the average over $\{Q_{j},P_{j}\}$.
@@
$H_{0}=\frac{1}{2}\sum_{j=1}^{N}\left\{P_{j}^{2}+\Omega_{j}^{2}\left[Q_{j}-\frac{c_{j}}{\Omega_{j}^{2}}\,X\right]^{2}\right\}+\frac{1}{2}\,\kappa\,X^{2},$ (109)
@@
Thus, the bath formed by the infinite Klein-Gordon chain is non-passive; it acts on the system with the systematic average force $\langle F\rangle=-\kappa\,X$. The effective stiffness $\kappa$ of that force as a function of $a=\omega_{a}/\omega_{0}$ is given by Eq. (110). It monotonically increases from 0 at $a=0$ to $k_{0}$ for $a\gg 1$. Both limiting values are intuitively plausible. For $a=0$, the model is equivalent to the Rubin model Zwanzig ; Weiss where the chain is not anchored and in the limit $N\gg 1$ (but not for finite $N$) serves as a passive bath, which does not produce a systematic force on the fixed system. For $a\gg 1$ the chain acts as an infinitely heavy anchor. In that case, the system feels only the spring of stiffness $k_{0}$ by which it is attached to the terminal atom of the chain.
@@ VII Klein-Gordon chain: Langevin equation
$F=-\frac{\partial H_{0}}{\partial X}=-\frac{\partial H_{0}^{\prime\prime}}{\partial X}=-k_{0}X+\sum_{j}c_{j}Q_{j},$ (116)
where $c_{j}$ are given by Eq. (103). Then $F_{0}(t)=e^{L_{0}t}F$ has the form
$F_{0}(t)=-k_{0}X+\sum_{j}c_{j}Q_{j}(t).$ (117)
The equations of motion for $Q_{j}(t)$, governed by Hamiltonian $H_{0}$, Eq. (106), read
@@
We observe that $F_{0}(t)$ satisfies conditions (87)-(89), mentioned at the end of Section 5, which ensure that statistical properties of the zero-centered noise $E_{0}(t)=F_{0}(t)-\langle F\rangle$ do not depend on $X$. We can also show that directly. Indeed, using the substitution
@@
stiffness $\kappa$ of the bath-induced potential is given by Eq. (110), and the dissipative kernel $\Gamma(t)$ is related to the zero-centered noise $E_{0}(t)$ by the standard fluctuation-dissipation relation
@@
and restricting to the leading approximation $F^{\dagger}(t)\approx E_{0}(t)$. One can show that for the Klein-Gordon model the next term $E_{1}(t)$, which is given by Eq. (78), vanishes. Indeed, as follows from Eqs. (122) and (126), $E_{0}(t)$ is linear in $X$, the derivative
@@
We finish this section by deriving a more explicit expression for the correlation $\langle EE_{0}(t)\rangle$ and the dissipation kernel $\Gamma(t)$. According to Eqs. (127) and (130), the kernel equals
@@
and the noise correlation equals $\langle EE_{0}(t)\rangle=\frac{M}{\beta}\,\Gamma(t)$. For the non-anchored chain ($a=0$), the dimensionless kernel $\gamma(t)$ is expressed in terms of the Bessel function $J_{1}(x)$,
$\gamma(t)=\frac{1}{2\,\omega_{0}\,t}\,J_{1}(\omega_{0}t).$ (137)
@@ VIII Conclusion
The concept of the PMF naturally emerges in equilibrium statistical mechanics of complex systems consisting of many particles. The force $F_{i}$ on $i$-th particle of the system is produced by the bath disturbed by the interaction with other parts of the system. As a result, even if the bath is a translationally-invariant fluid, the mean force $\langle F_{i}\rangle$ is generally not zero and is determined by the bath-induced part of the PMF, denoted above as $\Delta V$. With the PMF it is straightforward to evaluate average values in equilibrium, but there is a variety of views on how to incorporate the PMF in the Langevin dynamics.
In this paper, we have simplified the problem by assuming that the system of interest consists of a single particle. In that case, the PMF is usually irrelevant because the homogeneous bath is passive and produces no systematic force on a fixed system. That is not so, however, if the bath is not translationally-invariant. In the presented model the symmetry of the bath is broken by on-site anchor potentials applied to individual particles of the bath. Although that setting offers a simpler exposition, we were unable to derive a Langevin equation in a closed form. The non-passivity of the bath not only modifies the average potential acting on the system, but also makes the dissipation kernel and noise statistics depend on the position of the system $X$. That dependence is a priory unknown and there seems to be no general relation connecting it to the PMF. Therefore, we argue that there is no general form for a Langevin equation with the PMF. The conclusion corroborates the findings of Refs. Schilling ; Schilling_review ; Vroy , although we used here a different projection operator technique.
@@
$\dot{P}(t)=-\frac{\partial V_{ex}}{\partial X}-\kappa\,X(t)-\int_{0}^{t}\Gamma(t-\tau)\,P(\tau)\,d\tau+\zeta(t)-M\,X\,\Gamma(t),$ (A8)
@@
$\Gamma(t)=\frac{1}{M}\,\sum_{j=1}^{N}\left(\frac{c_{j}}{\Omega_{j}}\right)^{2}\cos(\Omega_{j}t),\quad\zeta(t)=\sum_{j=1}^{N}c_{j}Q_{j}^{0}(t),$ (A9)
@@
and $X=X(0)$. In the limit $N\to\infty$ $\kappa$ is given by expression (110). Finally, absorbing the term $M\,X\,\Gamma(t)$ into noise and using notations of the main text, we obtain